# An efficient Monte Carlo method for valid prior-free possibilistic statistical inference

### v2

@@ 6 Conclusion
The key innovation, motivated in part by Jiang et al., (2023), is to stitch together a collection of simple but individually-inadequate approximations into a superior meta-approximation. Here, for the individual approximations, I’m using the variational-like strategy put forward recently in Cella and Martin, (2024), rather than the bootstrap-based strategy offered in Jiang et al., (2023), because the former is computationally simpler than the latter. There may, however, be applications in which the additional flexibility offered by the bootstrap-based strategy is worth the extra computational costs.
The results here mark a first and important step in a series of developments leading to an IM toolbox for more-or-less off-the-shelf use by practitioners. The obvious next step is the development of general-purpose software for those common models—like the ones considered as examples in Sections 4 and 5 above—often encountered in applications. A natural next step would be extending the methods proposed here in various directions. This includes extensions to general non- and semiparametric models, where there may not be a likelihood function (e.g., Cella and Martin, 2022), which is well within reach since the application in Section 5 involved a semiparametric model. This would also include problems that involve “model uncertainty,” i.e., where all or part of what’s unknown and to be learned is the underlying model structure. A good example of this is mixture models, where the number of mixture components is unknown. Structure learning problems often require regularization, since the data alone cannot rule out overly-complex models, so this leads naturally into extensions of the proposed computational framework to accommodate cases with what Martin, 2022a ; Martin, 2022b has referred to as “partial prior information.” Finally, while the new proposal here, and the ideas in Cella and Martin, (2024) and elsewhere that it’s built on, aren’t specific to models involving low-dimensional unknowns, there are sure to be challenges associated with scaling up this proposal to handle high-dimensional problems. My current suggestion is to focus this effort on improving the efficiency and/or flexibility of the variational approximation in Section 3.3.
@@ Acknowledgments
This work is partially supported by the U.S. National Science Foundation, under grants SES–2051225 and DMS–2412628.
Appendix A Technical details
A.1 Proof of Theorem 1
@@
where the last inequality follows by the stochastically-no-smaller-than-${\sf Unif}(0,1)$ property of $\mathsf{M}$. To prove necessity, take the given $\mathsf{Q}\in\mathscr{C}(\overline{\Pi})$ and consider a random element $\Theta\sim\mathsf{Q}$. Note that $\Theta\in\partial C_{\pi(\Theta)}$ with $\mathsf{Q}$-probability 1. For some intuition, imagine partitioning $\mathbb{T}$ based on these level sets; then $\Theta$ itself is determined by the level set it’s on together with its position on the level set, and hence $\mathsf{Q}$ corresponds to an average of the conditional distribution of $\Theta$, given $\pi(\Theta)$, with respect to the marginal distribution of $\pi(\Theta)$. It follows from (8) that the random variable $\pi(\Theta)$ is stochastically no smaller than ${\sf Unif}(0,1)$, and let $\mathsf{M}$ be its marginal distribution relative to $\mathsf{Q}$. Similarly, take $\mathsf{K}^{\beta}$ to be (a version of) the conditional distribution of $\Theta$, given $\pi(\Theta)=\beta$, relative to $\mathsf{Q}$—note that $\mathsf{K}^{\beta}$ is fully supported on $\partial C_{\beta}\subset C_{\beta}$, as required. Then the equality (9) follows from the law of iterated expectation, completing the proof.
A.2 IMs and variational approximations
@@
The goal here is to review some of the relevant details presented in Cella and Martin, (2024) concerning the choice of $\xi=\xi(z,\alpha)$ employed in Section 3.3 above. The focus here, as above, is on the case where the variational family is Gaussian, but similar things can surely be done with other distributional families. The key features of the Gaussian are that its credible sets can be described in closed-form (ellipsoids) and that it’s straightforward to sample on the boundary of these credible sets.
Recall the Gaussian family of approximations $\mathsf{R}_{z}^{\xi}$ in (14), where $\xi$ is a $d$-vector with positive entries and $J_{z}(\xi)$ is the $\xi$-modified version of the observed Fisher information matirx $J_{z}$ with spectral decomposition $J_{z}=E\Lambda E^{\top}$. This is a generalization of a simpler but less flexible approximation that works with a scalar $\xi$ and sets $J_{z}(\xi)=\xi^{-2}\,J_{z}$. Also recall the Gaussian family’s $\alpha$-cut in (15):
$C_{\alpha}^{\xi}(z)=\{\theta:(\theta-\hat{\theta}_{z})^{\top}\,J_{z}(\xi)\,(\theta-\hat{\theta}_{z})\leq F_{d}^{-1}(1-\alpha)\},\quad\alpha\in[0,1].$
The key observation is that the Gaussian approximation $\mathsf{R}_{z}^{\xi}$ assigns probability at least $1-\alpha$ to the original IM’s $\alpha$-cut $C_{\alpha}(z)$ if $C_{\alpha}^{\xi}(z)\supseteq C_{\alpha}(z)$ or, equivalently, if
@@
Since the contour $\pi_{z}$ is itself approximately Gaussian (Martin and Williams, 2024) and the maximum likelihood estimator $\hat{\theta}_{z}$, also the mode of $\pi_{z}$, is in $C_{\alpha}^{\xi}(z)$, it follows that the action in the above supremum takes place on the boundary $\partial C_{\alpha}^{\xi}(z)$. Moreover, since equality in the above display implies a near-perfect match between the IM’s and the posited Gaussian $\alpha$-cuts, a reasonable goal is to find a root to the function
$g_{\alpha}(\xi):=\max_{\theta\in\partial C_{\alpha}^{\xi}(z)}\pi_{z}(\theta)-\alpha.$
Design of an iterative algorithm to find this root requires care, primarily because evaluating $\pi_{z}$ is expensive; so the goal is to evaluate $g_{\alpha}(\xi)$ with as few $\pi_{z}$ evaluations as possible. Cella and Martin, (2024) propose to represent the boundary of $C_{\alpha}^{\xi}(z)$ by $2d$-many vectors
$\vartheta_{s}^{\xi,\pm}:=\hat{\theta}_{n}\pm\bigl{\{}F_{d}^{-1}(1-\alpha)\,\xi_{s}\,/\,\lambda_{s}\}^{1/2}\,e_{s},\quad s=1,\ldots,d,$ (19)
@@
$\hat{g}_{\alpha,s}(\xi)=\max\{\pi_{z}(\vartheta_{s}^{\xi,+}),\pi_{z}(\vartheta_{s}^{\xi,-})\}-\alpha,\quad s=1,\ldots,d.$ (20)
At least intuitively, negative and positive $\hat{g}_{\alpha,s}(\xi)$ indicate that the $\alpha$-cut $C_{\alpha}^{\xi}(z)$ is too large and too small, respectively in the $e_{s}$-direction. From here, Cella and Martin, (2024) suggest applying a stochastic approximation algorithm à la robbinsmonro and kushner to construct a sequence $(\xi^{(t)}:t\geq 1)$ of $d$-vectors that converges to a root of $\hat{g}_{\alpha}$ and, hence, an approximate root of $g_{\alpha}$. For an initial guess $\xi^{(0)}$, the specific sequence is defined as
$\xi_{s}^{(t+1)}=\xi_{s}^{(t)}+w_{t+1}\,\hat{g}_{\alpha,s}(\xi^{(t)}),\quad s=1,\ldots,d,\quad t\geq 0,$
@@
These steps are iterated until (practical) convergence is achieved, and the limit is what I called $\xi(z,\alpha)$ in Section 3.3. The basic steps are outlined in Algorithm 1, but I refer to Cella and Martin, (2024) for further details and discussion.
@@
initialize: $\alpha$-level, guess $\xi^{(0)}$, step size sequence $(w_{t})$, and threshold $\varepsilon>0$;
@@
       construct the representative points $\{\vartheta_{s}^{\xi^{(t)},\pm}:s=1,\ldots,d\}$ as in (19);
       evaluate $\hat{g}_{\alpha,s}(\xi_{s}^{(t)})$ for $s=1,\ldots,d$ as in (20);
       update $\xi_{s}^{(t+1)}=\xi_{s}^{(t)}\pm w_{t+1}\,\hat{g}_{\alpha,s}(\xi_{s}^{(t)})$ for $s=1,\ldots,d$;
       if $\max_{s}|\xi_{s}^{(t+1)}-\xi_{s}^{(t)}|<\varepsilon$ then
             $\xi(z,\alpha)=\xi^{(t+1)}$;

### v3

@@ 7 Conclusion
The key innovation is to stitch together a collection of simple but individually-inadequate approximations into a superior meta-approximation. Here, for the individual approximations, I’m using the strategy advanced in Cella and Martin, (2025), rather than the bootstrap-based strategy offered in Jiang et al., (2024), because the former is computationally simpler than the latter. There may, however, be applications in which the additional flexibility offered by a bootstrap-based strategy is worth the extra computational costs. This question/topic will be investigated elsewhere.
The results here mark a first and important step in a series of developments leading to an IM toolbox for more-or-less off-the-shelf use by practitioners. The obvious next step is the development of general-purpose software for those common models—like the ones in the above examples—often encountered in applications. A natural next step would be extending the methods proposed here in various directions. This includes extensions to general non- and semiparametric models, where there may not be a likelihood function (e.g., Cella and Martin, 2022a ), which is well within reach. This would also include problems that involve “model uncertainty,” i.e., where all or part of what’s unknown and to be learned is the underlying model structure. A good example of this is mixture models, where the number of mixture components is unknown. Structure learning problems often require regularization, since the data alone cannot rule out overly-complex models, so this leads naturally into extensions of the proposed computational framework to accommodate cases with what Martin, 2022a ; Martin, 2022b has referred to as “partial prior information.” Finally, while the new proposal here, and the ideas in Cella and Martin, (2025) and elsewhere that it’s built on, aren’t specific to models involving low-dimensional unknowns, there are sure to be challenges associated with scaling up this proposal to handle high-dimensional problems. My current suggestion is to focus this effort on improving the efficiency and/or flexibility of the variational approximation in Section 3.3.
@@ Acknowledgments
This work is partially supported by the U.S. National Science Foundation, under grants SES–2051225 and DMS–2412628. Thanks to the anonymous reviewers for their feedback on an earlier version, and to Emily Hector, Chuanhai Liu, Sahil Patel, and James Robertson for helpful conversations and some valuable assistance.
Appendix A Possibility theory primer
Possibility measures (e.g., Dubois and Prade, 1988; Dubois, 2006) are among the simplest forms of imprecise probability, closely linked to fuzzy set theory (e.g., Zadeh, 1978) and Dempster–Shafer theory (e.g., Shafer, 1976, 1987). Probability and possibility theory differ philosophically and mathematically, but here I’ll only briefly discuss the latter.
The mathematical differences between probability and possibility theory can be succinctly summarized as follows:
Optimization is to possibility theory what integration is to probability theory.
That is, a possibility measure $\Pi$ defined on a space $\mathbb{Z}$ is determined by a function $\pi:\mathbb{Z}\to[0,1]$ with the property that $\sup_{z\in\mathbb{Z}}\pi(z)=1$. This function is called the possibility contour and the supremum-equals-1 property is a normalization condition akin to the integral-equals-1 property satisfied by probability densities. Then, naturally, the possibility measure $\Pi$ is determined by optimizing its contour, i.e., $\Pi(A)=\sup_{z\in A}\pi(z)$, for any $A\subseteq\mathbb{Z}$, just like a probability measure is determined by integrating its density.
The possibility calculus just described is conceptually straightforward and easy to analyze but, of course, there may be computational challenges associated with implementing the theory—just like with probability. Moreover, this optimization-centric framework of uncertainty quantification is no less rigorous than the familiar probability theory based on (Lebesgue) integration. To justify this latter claim, I’ll make the following two points.
First, and of particular relevance for the developments in this paper, the aforementioned supremum-equals-1 normalization condition ensures that $\Pi$ is a coherent upper probability (e.g., De Cooman, 1997; De Cooman and Aeyels, 1999; Walley, 1997) in the spirit of Walley, (1991) and others. Among other things, this means that $\Pi$ determines a non-empty (closed and convex) set of ordinary probabilities that it dominates:
$\mathscr{C}(\Pi)=\{\mathsf{Q}\in\text{probs}(\mathbb{Z}):\mathsf{Q}(H)\leq\Pi(H)\text{ for all measurable $H$}\},$ (16)
where $\text{probs}(\mathbb{Z})$ is the set of probabilities supported on the Borel $\sigma$-algebra of measurable subsets in $\mathbb{Z}$. The set $\mathscr{C}(\Pi)$ is called the credal set and all (coherent) upper probabilities have one. Aside from being relatively simple, a key advantage to possibilistic uncertainty quantification is that the associated credal set has a statistically oriented characterization, which I discuss in the main paper.
Second, one can extend possibilistic evaluations beyond just the possibility values assigned to hypotheses. This extension operation is analogous to the evaluation of expected values via Lebesgue integration of measurable functions. This extension relies on what’s called Choquet integration, originally developed in Choquet, (1954), thoroughly analyzed in Troffaes and de Cooman, (2014) and elsewhere, with applications in statistics (Huber, 1973, e.g.,). Note that Choquet integration is not specifically tied to possibility theory, it’s for integration with respect to general non-additive capacities, and possibility measures are a special case; there are, however, some nice simplifications that happen because of the maxitive—as opposed to additive—form of possibility measures. Suppose that $h:\mathbb{T}\to\mathbb{R}$ is a non-negative function; non-negativity is just for convenience here. Then the Choquet integral of $h$ with respect to the possibility measure $\Pi$ is given by
[…]
where $\pi$ is the contour function associated with $\Pi$. As an important special case, connecting the possibilistic marginalization strategies discussed in Appendix F, suppose that $h(\theta)=1(\theta\in H)$, where $H=\{\theta:k(\theta)\in K\}$ for some function $k$ and some subset $K\subseteq k(\mathbb{T})$. Then applying the Choquet integral formula in (17) gives
$\Pi\,h=\sup_{\theta:h(\theta)=1}\pi(\theta)=\sup_{\theta\in H}\pi(\theta)=\sup_{\kappa\in K}\sup_{\theta:k(\theta)=\kappa}\pi(\theta).$
There’s a clear analogy that can be made between the right-hand side of the above display and the more familiar Bayesian calculus: the degree of possibility/probability assigned to “$\Theta\in H$” or, equivalently, to “$k(\Theta)\in K$” is obtained by first getting the marginal contour/density for $k(\Theta)$ at $\kappa$ by optimizing/integrating over $\{\theta:k(\theta)=\kappa\}$ and then optimizing/integrating over $\kappa\in K$. The same can be said for Choquet integrals of other kinds of functions $h$, but the point is most easily made when $h$ is an indicator.
Appendix B Proof of Theorem 1 and Corollary 1
@@
where the last inequality follows by the stochastically-no-smaller-than-${\sf Unif}(0,1)$ property of $\mathsf{M}$. To prove necessity, take the given $\mathsf{Q}\in\mathscr{C}(\Pi)$ and consider a random element $\Theta\sim\mathsf{Q}$. Note that $\Theta\in\partial C_{\pi(\Theta)}$ with $\mathsf{Q}$-probability 1. For some intuition, imagine partitioning $\mathbb{T}$ based on these level sets; then $\Theta$ itself is determined by the level set it’s on together with its position on the level set, and hence $\mathsf{Q}$ corresponds to an average of the conditional distribution of $\Theta$, given $\pi(\Theta)$, with respect to the marginal distribution of $\pi(\Theta)$. It follows from (7) that the random variable $\pi(\Theta)$ is stochastically no smaller than ${\sf Unif}(0,1)$, and let $\mathsf{M}$ be its marginal distribution relative to $\mathsf{Q}$. Similarly, take $\mathsf{K}^{\beta}$ to be (a version of) the conditional distribution of $\Theta$, given $\pi(\Theta)=\beta$, relative to $\mathsf{Q}$—note that $\mathsf{K}^{\beta}$ is fully supported on $\partial C_{\beta}\subset C_{\beta}$, as required. Then the equality (8) follows from the law of iterated expectation, completing the proof of Theorem 1.
Equality throughout the above display is achieved if, first, $\mathsf{K}^{\beta}(C_{\alpha})=0$ for all $\alpha<\beta$ and, second, if $\mathsf{M}$ is ${\sf Unif}(0,1)$. This condition on the kernel is achieved if $\mathsf{K}^{\beta}$ is supported on the boundary, $\partial C_{\beta}$, of $C_{\beta}$, for each $\beta\in[0,1]$, hence Corollary 1.
Appendix C Direct implementation of Corollary 1
Corollary 1 describes the inner probabilistic approximation of the IM $\Pi_{z}$ as a mixture of kernels $\mathsf{K}_{z}^{\alpha}$ supported on the boundary $\partial C_{\alpha}(z)$ of the possibility contour’s $\alpha$-cut, for each $\alpha\in[0,1]$. The reader might mistakingly think that, because $\mathsf{K}_{z}^{\alpha}$ is supported on a lower-dimensional space, handling such a mixture is an impractically impossible task. The cause of this confusion is that the reader imagines having to condition on the probability-0 event associated with the lower-dimensional boundary. While conditioning on the boundary event is one way to construct such a kernel, that’s not the only way. Instead, I can just define $\mathsf{K}_{z}^{\alpha}$ to be a distribution supported on $\partial C_{\alpha}(z)$, and there would rarely be problems with carrying this out. For instance, an incredibly mild assumption is that each $C_{\alpha}(z)$ be star-convex, meaning that there exists a point on the interior (e.g., the maximum likelihood estimator $\hat{\theta}_{z}$) such that every point on the boundary can be uniquely identified by a ray emanating from that interior point. The cartoon illustration in Figure 1 in the main text is star-convex but not elliptical/convex. With this structure, sampling on the boundary $\partial C_{\alpha}(z)$ is equivalent to sampling on the unit ball, which is not particularly challenging. Then one just needs to identify points on the unit ball with their corresponding points on $\partial C_{\alpha}(z)$, which would generally require some iterative numerical methods; see below. An advantage of the second-layer approximation with an ellipsoidal $C_{\alpha}^{\text{out}}(z)$ is that the mapping from the ball to $\partial C_{\alpha}^{\text{out}}(z)$ is explicit.
The main text made the claim that the elliptically-shaped contour formulation is not crucial the development; such an assumption is imposed solely because (a) it’s the “right” choice for large-samples so there’s little room for improvement, (b) it makes the computations relatively simple, and (c) there is no empirical evidence indicating that this approximation is inaccurate. I make no claims, however, that the elliptical formulation is universal. Fortunately, it’s easy to extend the method to handle other forms, even the inner probabilistic approximation exactly. Below I’ll describe implementation of the inner probabilistic approximation itself since that’s arguably the most general and the most complicated. Other problem-specific approximations could be considered too, e.g., with $C_{\alpha}^{\sigma}(z)$ a suitable Kullback–Leibler neighborhood when the parameter space is the simplex, but these details are beyond the scope of the present paper.
For this discussion, suppose that the level sets of $\pi_{z}$ are non-elliptical (like in Figure 1 in the main text) but are star-convex in the sense described above. This is a very mild supposition. Then to sample on the boundary $C_{\alpha}(z)$, I just need to sample on the unit ball, i.e., sample a random direction $U$ in $d$-space, and then follow the line in that direction until it intersects with $\partial C_{\alpha}(z)$. That is, solve the equation $\pi_{z}(\hat{\theta}_{z}+tU)=\alpha$ as a function of $t>0$. Since $\pi_{z}$ can only be evaluated using Monte Carlo, this root-solving problem can’t be solved using Newton’s method or bisection—it requires something like the stochastic approximation algorithm employed by Cella and Martin as discussed in Appendix D. To summarize, independent samples from the inner probabilistic approximation $\mathsf{Q}_{z}^{\text{inn}}$ are drawn by iterating the following two steps:
  1. 1.
    Draw $A\sim{\sf Unif}(0,1)$.
  2. 2.
    Given $A=\alpha$, draw $U$ uniformly on the unit $d$-dimensional ball (e.g., by drawing a $d$-dimensional standard Gaussian and then normalizing it) and then set $\Theta$ equal to the root of the function $g(t)=\pi_{z}(\hat{\theta}_{z}+tU)-\alpha$, $t>0$.
While carrying out the above computations are not difficult, they’re more expensive than the solution advocated for in the main text based on the elliptical approximation. At least part of the reason for this increased expense is that one needs to perform the root-solving $M$ many times, where $M$ is the desired Monte Carlo sample size, typically in the thousands, whereas the strategy described in the paper only needs root-solving to find $\sigma(z,\alpha)$ on a fixed set of about 100 different $\alpha$ values; see Remark 1 below.
Again, no claims are made that the elliptical approximation is universal, but it’s simple, asymptotically optimal, and works well in a wide range of examples, including those with relatively small sample sizes. Future investigations will explore IM solutions to more complex problems where the full generality of the proposed computational strategy, as described above, will surely be needed.
Appendix D Cella and Martin’s algorithm
@@
The goal here is to review some of the relevant details presented in Cella and Martin, (2025) concerning the choice of $\sigma=\sigma(z,\alpha)$ employed in Section 3.3 in the main text. The focus here is on the case where the $\alpha$-cuts $C_{\alpha}^{\sigma}(z)$ are ellipsoidal, but similar things can be done with other shapes, as discussed in Appendix C above.
Recall the eigenvalue-adjusted version $J_{z}(\sigma)$ of the observed Fisher information matrix $J_{z}$, i.e., based on the spectral decomposition $J_{z}=E\Lambda E^{\top}$. Also recall the Gaussian possibility’s $\alpha$-cut in (12):
$C_{\alpha}^{\sigma}(z)=\{\theta:(\theta-\hat{\theta}_{z})^{\top}\,J_{z}(\sigma)\,(\theta-\hat{\theta}_{z})\leq F_{d}^{-1}(1-\alpha)\},\quad\alpha\in[0,1].$
The key observation is that the desired containment $C_{\alpha}^{\sigma}(z)\supseteq C_{\alpha}(z)$ holds if and only if
@@
Since the contour $\pi_{z}$ is itself approximately Gaussian (Martin and Williams, 2025) and the maximum likelihood estimator $\hat{\theta}_{z}$, also the mode of $\pi_{z}$, is in $C_{\alpha}^{\sigma}(z)$, it’s expected that the above supremum is attained on the boundary $\partial C_{\alpha}^{\sigma}(z)$. Moreover, since equality in the above display implies a near-perfect match between the IM’s and proposed elliptical $\alpha$-cuts, a reasonable goal is to find a root to the function
$g_{\alpha}(\sigma):=\max_{\theta\in\partial C_{\alpha}^{\sigma}(z)}\pi_{z}(\theta)-\alpha.$
Design of an iterative algorithm to find this root requires care, primarily because evaluating $\pi_{z}$ is expensive; so the goal is to evaluate $g_{\alpha}(\sigma)$ with as few $\pi_{z}$ evaluations as possible. Cella and Martin, (2025) propose to represent the boundary of $C_{\alpha}^{\sigma}(z)$ by $2d$-many vectors
$\vartheta_{s}^{\sigma,\pm}:=\hat{\theta}_{z}\pm\bigl{\{}F_{d}^{-1}(1-\alpha)\,\sigma_{s}\,/\,\lambda_{s}\}^{1/2}\,e_{s},\quad s=1,\ldots,d,$ (18)
@@
$\hat{g}_{\alpha,s}(\sigma)=\max\{\pi_{z}(\vartheta_{s}^{\sigma,+}),\pi_{z}(\vartheta_{s}^{\sigma,-})\}-\alpha,\quad s=1,\ldots,d.$ (19)
At least intuitively, negative and positive $\hat{g}_{\alpha,s}(\sigma)$ indicate that the $\alpha$-cut $C_{\alpha}^{\sigma}(z)$ is too large and too small, respectively in the $e_{s}$-direction. From here, Cella and Martin, (2025) suggest applying a stochastic approximation algorithm à la Robbins and Monro, (1951) and Kushner and Yin, (2003) to construct a sequence $(\sigma^{(t)}:t\geq 1)$ of $d$-vectors that converges to a root of $\hat{g}_{\alpha}$ and, hence, an approximate root of $g_{\alpha}$. For an initial guess $\sigma^{(0)}$, the specific sequence is defined as
$\sigma_{s}^{(t+1)}=\sigma_{s}^{(t)}+w_{t+1}\,\hat{g}_{\alpha,s}(\sigma^{(t)}),\quad s=1,\ldots,d,\quad t\geq 0,$
@@
These steps are iterated until (practical) convergence is achieved, and the limit is what I called $\sigma(z,\alpha)$ in Section 3.3. The basic steps are outlined in Algorithm 2, but I refer to Cella and Martin, (2025) for further details and discussion.
@@
initialize: $\alpha$-level, guess $\sigma^{(0)}$, step size sequence $(w_{t})$, and threshold $\varepsilon>0$;
@@
       construct the representative points $\{\vartheta_{s}^{\sigma^{(t)},\pm}:s=1,\ldots,d\}$ as in (18);
       evaluate $\hat{g}_{\alpha,s}(\sigma_{s}^{(t)})$ for $s=1,\ldots,d$ as in (19);
       update $\sigma_{s}^{(t+1)}=\sigma_{s}^{(t)}\pm w_{t+1}\,\hat{g}_{\alpha,s}(\sigma_{s}^{(t)})$ for $s=1,\ldots,d$;
       if $\max_{s}|\sigma_{s}^{(t+1)}-\sigma_{s}^{(t)}|<\varepsilon$ then
             $\sigma(z,\alpha)=\sigma^{(t+1)}$;